# Casimir interactions and drift currents

### v1

@@ I Introduction
Quantum vacuum fluctuations of electromagnetic fields give rise to ubiquitous coupling between objects regardless of their properties. The associated Casimir force [Casimir1948], in which the finite speed of light $c$ must be taken in the electromagnetic exchange, has been of great interest. It probes the interplay between the fundamental properties and geometry of the interaction objects, but it is also important for nano and micromachines for their sticktion and adhesion phenomena [Woods2016, Klimchitskaya]. The discoveries of new materials have stimulated the Casimir field since properties of novel Dirac materials, topological insulators, and quasi-one dimensional systems have led to a diverse set of scaling laws, quantization effects, and even repulsion [Woods2017, Rodriguez2014, Nefedov2021]. Casimir interactions under non-equilibrium conditions have also been of much interest recently. For example, holding the objects at different temperatures or moving one object with respect to the other have been studied by different groups [Bimonte2017, Antezza2008, Volokitin2011, Tomassone1997, Intravaia2014]. In addition, theoretical studies have shown that when two charge-neutral objects move parallel to each other, quantum electromagnetic fluctuations would produce a drag force even in vacuum [Milton2016, Milton2023].
@@
Casimir phenomena are also influenced by having drift charge carriers, for example in the case of bias voltage resulting in dc currents in the objects. Drift carriers affect the electromagnetic waves inside each material, and they also couple to the vacuum fluctuations exchanged between the objects. The case of semiconducting plates supporting dc currents have been considered where the dielectric response was taken via the semi-classical Boltzmann equation [Dalvit2008] or the Drude model [Shapiro2010]. The Casimir-Polder interaction between a nanoparticle and a current-carrying substrate was also studied in [Shapiro], where one distinguishes between two models for the dielectric response of the substrate. In one model it is assumed that the fluctuations arise predominantly from the lattice of the material taken to be in equilibrium, while the other model takes the electron plasma approximation with Doppler–shifted noise as the dominant source with an idealized separation of contributions from lattice and electrons [Shapiro].
@@
In this paper, we examine the Casimir interaction between two dc current carrying parallel graphene sheets. As found in the previous papers [Volokitin_2013, Farias], stronger effects are found in systems with larger currents. Graphene can support carriers with very large drift velocities [Yamoah2017, Yin2014, Dorgan2010], thus this effect may be stronger in interacting graphene sheets. For this purpose, we resolve the boundary condition of two parallel graphene sheets carrying a drift current. Then the Maxwell stress tensor is used to find the fluctuation induced forces in terms of the reflection coefficients of the two layers.
@@
Similar setup has been considered by [Volokitin_2013] where the effect of the drift current is treated simply as a Doppler shift. Instead in our paper we utilize a phenomenological shifted Fermi disk (SFD) model for the conductivity, which more accurately captures the dielectric response of the current carrying graphene and reduces to a Doppler-like transformation only for two dimensional electron gas [Sabbaghi1, Sabbaghi2].
@@
Unlike earlier approaches that represent the effect of a drift current as a simple Doppler shift in the electromagnetic response or through semiclassical Drude–Boltzmann transport [Shapiro, Shapiro2010, Volokitin_2013], the SFD model also captures the intrinsic redistribution of carriers in momentum space caused by steady-state drift. This microscopic description naturally accounts for the graphene’s linear Dirac dispersion and the resulting anisotropic conductivity tensor, thereby going beyond the Doppler-like approximations valid only for parabolic two-dimensional electron gases [Sabbaghi1, Sabbaghi2]. The SFD model thus provides a physically consistent framework for treating non-equilibrium optical response in current-carrying graphene.
@@ II Graphene layers and current induced modifications in the optical response
An integral part of our modeling is the graphene optical response in the presence of a drift current. The preferred direction established by the current breaks the local rotational invariance and the time reversal symmetry in the nonlocal response leading to modified plasmonic group velocities [Sabbaghi1, Sabbaghi2]. In the presence of a steady-state drift current, the electrons reach a non-equilibrium satationary state by acquiring momentum and kinetic energy from the drain-source voltage while simultaneously dissipating part of that energy through electron scattering mechanisms.
@@ III The Maxwell stress tensor and force components
The fluctuation induced force of the electromagnetic excitations is closely related to the Maxwell stress tensor $\stackrel{{\scriptstyle\leftrightarrow}}{{\mathscr{T}}}$ [Lifshitz, Polder]. In addition to the interaction directed along the $z$-axis, the drift current induces a lateral Casimir force component. Considering the case of currents flowing in the $x$-direction, these are given as
@@
where $E_{x,y,z}$ and $B_{x,y,z}$ are the components of the electric and magnetic fields, respectively. All $\langle{...}\rangle$ terms correspond to fluctuations of the fields, expressed through their correlation functions evaluated at $z=0_{\pm}$, i.e., as the boundary is approached from the top and bottom of the $z$-axis, respectively (see Fig. 1). In alternative formulations, these correlations are often obtained from the electromagnetic Green’s functions of the system [Volokitinbook]. In that approach, the fluctuating sources are encoded through the dyadic Green’s tensors, which propagate the fields subject to the given boundary conditions, thereby providing a natural link between material response and fluctuating fields. By contrast, in Rytov’s fluctuational electrodynamics [Lifshitz], the starting point is the correlation functions of the fluctuating current densities inside the media. The field correlations are then derived by propagating these stochastic sources through Maxwell’s equations. The Green’s function and Rytov approaches are formally equivalent, but differ in emphasis: the former focuses on constructing the field propagators explicitly, while the latter emphasizes the statistical properties of the microscopic sources. In this work we adopt the Rytov framework, which is particularly convenient for treating boundary conditions at planar interfaces and for expressing fluctuation–dissipation relations directly in terms of material response functions.
To calculate explicitly the force components in Eqs. (4)-(5), the electromagnetic boundary conditions are resolved for the planar symmetry in Fig. 1. Details can be found in Appendix A. The current fluctuations in the graphene layers are taken into account via the fluctuation dissipation theorem [Lifshitz, Polder], for which $\langle j_{f\alpha}(\bm{q},\omega)j_{f\beta}(\bm{q},\omega)\rangle=\frac{\hbar\omega}{\pi}\left(\frac{1}{2}+n_{B}(\omega)\right){\rm Re}[\sigma_{\alpha\beta}(\bm{q},\omega)]$ where $n_{B}(\omega)=1/(e^{(\hbar\omega)/k_{B}T}-1)$ with $T$ denoting temperature. In the case of drift carriers flowing through each graphene layer, this expression needs to be modified since the fluctuation dissipation theorem is valid in equilibrium. Thus, the original stationary reference frame is changed via a Lorentz transformation to a frame attached to the drifting carriers, which means that the conductivity must take into account the presence of current density with magnitude $I_{d}\approx n_{s}\beta_{d}v_{F}$ (model discussed earlier). Assuming the drift velocity is much smaller than the speed of light, one can omit $(v_{d}/c)^{2}$ and higher corrections in the Lorentz transformation [Volokitinbook]. As a result, there is a Doppler shifted frequency, $\omega^{\prime}=\omega-\bm{q}\cdot\bm{v}_{d,1}$ for the top layer and $\omega^{\prime\prime}=\omega-\bm{q}\cdot\bm{v}_{d,2}$ for the bottom layer.
@@
Eqs. (III)-(III) show that the vertical and lateral components of the current-modified Casimir interaction are composed of propagating ($q<\omega/c$) and evanescent ($q>\omega/c$) contributions, similar to the graphene-graphene interaction in equilibrium conditions [physics5040066]. We note that in the absence of a current, $F_{x}$ becomes zero as all the integrands are odd in $q_{x}$ leading to exact cancellations upon $q_{x}$ and $q_{x}$ integration. We also find that in this case, Eq. (III) can be evaluated by rotating the frequency integral into the complex plane, which transforms the $F_{z}$ expression into the conventional Lifhsitz-like form in the imaginary frequency domain [Woods2010, Klimchitskaya]. As a result, the Casimir force between the graphene sheets in the absence of drift currents is obtained as $F_{0}=-\frac{3\hbar c\alpha}{32\pi d^{4}}$ where $\alpha$ is the fine structure constant.
@@
At this point it is important to note that the Casimir interaction in the presence of drift currents has also been studied in a system composed of a nanoparticle and a 3D substrate, which is also relevant for the results obtained in Eqs. (III)-(III). In particular, Shapiro [Shapiro] developed a microscopic model for fluctuation-induced interactions between a nanoparticle and a current-carrying medium, identifying distinct physical pictures depending on whether the underlying fluctuations originate from the drifting electron plasma or from the lattice. In one of the considered scenarios, the contribution to the force comes from the moving electron plasma, while the lattice dissipation is neglected; in the other case, the fluctuation of the lattice dominates and thus the moving electrons would not affect the fluctuation-induced interaction. If the primary source of fluctuations originates from the lattice, it would be necessary to consider an effective temperature of the lattice that is different than the effective temperature of the drifting electrons, which complicates the problem as it is difficult to find these two different temperatures [Shapiro]. For the fluctuations originating primarily from the electrons a Doppler shift is induced in the fluctuation frequencies $\omega^{\prime},\omega^{\prime\prime}$ due to the carrier motion. In this work, we regard electrons as the primary source of fluctuations with a drift current in stationary medium.
Volokitin and Persson [Volokitin_2013] modeled the drift current in one of the two parallel graphene layers via a Lorentz transformation—formally equivalent to a current-free medium moving along the $x$-axis at $v_{d}$. However, this treatment differs fundamentally from the physical situation in graphene, where the lattice remains static in the laboratory frame while only the carriers are moving. In the electron rest frame induced by the Lorentz boost, the lattice appears to move oppositely, yielding a dielectric response that deviates from the equilibrium case of static electrons and lattice. Therefore, a simple Doppler shift cannot adequately describe the graphene conductivity under drift. In [Sabbaghi1] it has been shown that the SFD model and Doppler-shift approximation coincide only for an ideal two-dimensional electron gas, with discrepancies arising in graphene due to its Dirac nature. Here, we adopt a similar framework as in [Volokitin_2013] while implementing the full SFD model to achieve a comprehensive description of the graphene optical response under drift.
@@ IV Current induced modifications in the interaction
To understand the role of drift currents in the graphene Casimir interaction, here we present results for the numerical calculations of Eqs. (III)-(III) for the case of charge carriers flowing along the $x$-axis in the top graphene layer with Doppler shifted frequency $\omega^{\prime}=\omega-q_{x}v_{\rm d}$ (for the bottom layer there is no drift velocity). The results in Fig. 3 (a)-(b) show how the modified interaction along the vertical direction depends on $\beta_{d}$ and $d$. At $\beta_{d}=0$ when there is no current, $\Delta F_{z}=0$ and we recover the usual equilibrium Casimir force between two graphene sheets $F_{0}=-\frac{3\hbar c\alpha}{32\pi d^{4}}$ previously discussed in literature [Woods2017, Woods2010]. In the presence of a drift current, the Casimir force $F_{z}$ is modified by $\Delta F_{z}$ found to be with a positive sign. This repulsive graphene-graphene correction along the $z$-axis has a slightly slower than a quadratic $\beta_{d}^{2}$ dependence, as shown in Fig. 3 (a). We also note that $\Delta F_{z}$ has a different distance dependence when compared to $F_{0}$; in particular, we find that $\Delta F_{z}\sim\frac{1}{d^{5.8}}$ indicating a shorter range interaction compared to the equilibrium case with its $F_{0}\sim\frac{1}{d^{4}}$ distance power law. Fig. 3 (a, b) shows that this current-induced repulsive correction is most significant at small separations and large drift currents.
@@
As discussed earlier, due to the drifting electrons there is also a lateral Casimir-like force $F_{x}$ responsible for dissipating energy in the system. The lateral force opposes the direction of the driting carriers, and it is similar to the noncontact friction arising between two graphene sheets in relative motion [Shapiro, Farias, Gorbachev2012] . Fig. 3 (c) shows that the lateral force is several orders of magnitude smaller when compared to the equilibrium $F_{0}$. It has a linear dependence upon $\beta_{d}$ for $\beta_{d}<1$, which is consistent with previously reported results for small drift current [Volokitin_2013], and it has a nonlinear behavior as $\beta_{d}$ becomes larger. In Fig. 3 (d) we also show how $F_{x}$ changes as a function of the graphene-graphene separation. By interpolating the data, it is obtained that $F_{x}\sim\frac{1}{d^{4.2}}$ indicating that the lateral force has a somewhat shorter range coupling compared to $F_{0}$.
@@
It is also interesting to see how other factors, in addition to the interlayer distance and drift velocity, affect the lateral and vertical force components of the interaction. Fig. 4 (a) and (c) shows almost linear dependence upon temperature. The origin of this linear behavior can be traced to the thermal Bose factor entering the fluctuation-induced force. In the relevant low-frequency regime, where $\hbar\omega\ll k_{B}T$, the Bose–Einstein distribution admits the expansion $\left(\frac{1}{2}+n_{B}(\omega)\right)\approx\frac{k_{B}T}{\hbar\omega}+\mathcal{O}\left(\frac{\hbar\omega}{k_{B}T}\right)$, which shows that $\Delta F_{z}$ and $F_{x}$ become proportional to $k_{B}T/\hbar\omega$ to leading order. Another property that can also be used to modify the interaction is the graphene Fermi energy $E_{F}$. Increasing $E_{F}$ enhances the metallic-like nature of the materials leading to stronger Casimir interaction [Woods2010, Woods2017, Fialkovsky]. This is also observed in the case of current-modified graphene coupling. As shown in Fig. 4 (b) and (d), both the lateral and vertical forces experience direct correlation with $E_{F}$ with a close to linear functional behavior.
@@
At the end we discuss the case of currents flowing in both graphene layers but in the same direction. One notes that the two major factors in the Casimir energy in Eqs. (III)-(III) come from the Doppler shifted frequencies $\omega^{\prime}=\omega^{\prime\prime}=\omega-\bm{q}\cdot\bm{v}_{d}$ and the modified graphene conductivities according to the SFD model in Eq. (3). We can see that in this case the integrand in Eq. (III) is zero, which means that there is no lateral force. An analogous situation may also occur when the two graphene layers are physically moving with the same velocity in same directions [Volokitin_2001, Intravaia2014, Milton2016]. The non-contact frictional force in this case is zero since there is no relative motion between electrons in the two layers. The correction to the vertical interaction $\Delta F_{z}$ is also small compared to the results in Figs.3 and 5. Since the Doppler shifted frequencies are not relevant due to absence of relative motion between the electrons of the two layers, only the SFD modified conductivities are important, therefore the corresponding correction to the equilibrium vertical component is also less prominent.

### v2

# Casimir interactions between two parallel graphene sheets carrying steady-state drift currents

@@ I Introduction
Quantum vacuum fluctuations of electromagnetic fields give rise to ubiquitous coupling between objects regardless of their properties. The associated Casimir force [3], in which the finite speed of light $c$ must be taken in the electromagnetic exchange, has been of great interest. It probes the interplay between the fundamental properties and geometry of the interaction objects, but it is also important for nano and micromachines for their sticktion and adhesion phenomena [32, 15]. The discoveries of new materials have stimulated the Casimir field since properties of novel Dirac materials, topological insulators, and quasi-one dimensional systems have led to a diverse set of scaling laws, quantization effects, and even repulsion [22, 21, 19]. Casimir interactions under non-equilibrium conditions have also been of much interest recently. For example, holding the objects at different temperatures or moving one object with respect to the other have been studied by different groups [2, 1, 29, 27, 14]. In addition, theoretical studies have shown that when two charge-neutral objects move parallel to each other, quantum electromagnetic fluctuations would produce a drag force even in vacuum [18, 12].
@@
Casimir phenomena are also influenced by having drift charge carriers, for example in the case of bias voltage resulting in dc currents in the objects. Drift carriers affect the electromagnetic waves inside each material, and they also couple to the vacuum fluctuations exchanged between the objects. The case of semiconducting plates supporting dc currents have been considered where the dielectric response was taken via the semi-classical Boltzmann equation [4] or the Drude model [25]. The Casimir-Polder interaction between a nanoparticle and a current-carrying substrate was also studied in [26], where one distinguishes between two models for the dielectric response of the substrate. In one model it is assumed that the fluctuations arise predominantly from the lattice of the material taken to be in equilibrium, while the other model takes the electron plasma approximation with Doppler–shifted noise as the dominant source with an idealized separation of contributions from lattice and electrons [26].
@@
In this paper, we examine the Casimir interaction between two dc current carrying parallel graphene sheets. As found in the previous papers [30, 8], stronger effects are found in systems with larger currents. Graphene can support carriers with very large drift velocities [34, 35, 6], thus this effect may be stronger in interacting graphene sheets. For this purpose, we resolve the boundary condition of two parallel graphene sheets carrying a drift current. Then the Maxwell stress tensor is used to find the fluctuation induced forces in terms of the reflection coefficients of the two layers.
@@
Similar setup has been considered by [30] where the effect of the drift current is treated simply as a Doppler shift. Instead in our paper we utilize a phenomenological shifted Fermi disk (SFD) model for the conductivity, which more accurately captures the dielectric response of the current carrying graphene and reduces to a Doppler-like transformation only for two dimensional electron gas [23, 24].
@@
Unlike earlier approaches that represent the effect of a drift current as a simple Doppler shift in the electromagnetic response or through semiclassical Drude–Boltzmann transport [26, 25, 30], the SFD model also captures the intrinsic redistribution of carriers in momentum space caused by steady-state drift. This microscopic description naturally accounts for the graphene’s linear Dirac dispersion and the resulting anisotropic conductivity tensor, thereby going beyond the Doppler-like approximations valid only for parabolic two-dimensional electron gases [23, 24]. The SFD model thus provides a physically consistent framework for treating non-equilibrium optical response in current-carrying graphene.
@@ II Graphene layers and current induced modifications in the optical response
An integral part of our modeling is the graphene optical response in the presence of a drift current. The preferred direction established by the current breaks the local rotational invariance and the time reversal symmetry in the nonlocal response leading to modified plasmonic group velocities [23, 24]. In the presence of a steady-state drift current, the electrons reach a non-equilibrium satationary state by acquiring momentum and kinetic energy from the drain-source voltage while simultaneously dissipating part of that energy through electron scattering mechanisms.
@@ III The Maxwell stress tensor and force components
The fluctuation induced force of the electromagnetic excitations is closely related to the Maxwell stress tensor $\stackrel{{\scriptstyle\leftrightarrow}}{{\mathscr{T}}}$ [17, 20]. In addition to the interaction directed along the $z$-axis, the drift current induces a lateral Casimir force component. Considering the case of currents flowing in the $x$-direction, these are given as
@@
where $E_{x,y,z}$ and $B_{x,y,z}$ are the components of the electric and magnetic fields, respectively. All $\langle{...}\rangle$ terms correspond to fluctuations of the fields, expressed through their correlation functions evaluated at $z=0_{\pm}$, i.e., as the boundary is approached from the top and bottom of the $z$-axis, respectively (see Fig. 1). In alternative formulations, these correlations are often obtained from the electromagnetic Green’s functions of the system [31]. In that approach, the fluctuating sources are encoded through the dyadic Green’s tensors, which propagate the fields subject to the given boundary conditions, thereby providing a natural link between material response and fluctuating fields. By contrast, in Rytov’s fluctuational electrodynamics [17], the starting point is the correlation functions of the fluctuating current densities inside the media. The field correlations are then derived by propagating these stochastic sources through Maxwell’s equations. The Green’s function and Rytov approaches are formally equivalent, but differ in emphasis: the former focuses on constructing the field propagators explicitly, while the latter emphasizes the statistical properties of the microscopic sources. In this work we adopt the Rytov framework, which is particularly convenient for treating boundary conditions at planar interfaces and for expressing fluctuation–dissipation relations directly in terms of material response functions.
To calculate explicitly the force components in Eqs. (4)-(5), the electromagnetic boundary conditions are resolved for the planar symmetry in Fig. 1. Details can be found in Appendix A. The current fluctuations in the graphene layers are taken into account via the fluctuation dissipation theorem [17, 20], for which $\langle j_{f\alpha}(\bm{q},\omega)j_{f\beta}(\bm{q},\omega)\rangle=\frac{\hbar\omega}{\pi}\left(\frac{1}{2}+n_{B}(\omega)\right){\rm Re}[\sigma_{\alpha\beta}(\bm{q},\omega)]$ where $n_{B}(\omega)=1/(e^{(\hbar\omega)/k_{B}T}-1)$ with $T$ denoting temperature. In the case of drift carriers flowing through each graphene layer, this expression needs to be modified since the fluctuation dissipation theorem is valid in equilibrium. Thus, the original stationary reference frame is changed via a Lorentz transformation to a frame attached to the drifting carriers, which means that the conductivity must take into account the presence of current density with magnitude $I_{d}\approx n_{s}\beta_{d}v_{F}$ (model discussed earlier). Assuming the drift velocity is much smaller than the speed of light, one can omit $(v_{d}/c)^{2}$ and higher corrections in the Lorentz transformation [31]. As a result, there is a Doppler shifted frequency, $\omega^{\prime}=\omega-\bm{q}\cdot\bm{v}_{d,1}$ for the top layer and $\omega^{\prime\prime}=\omega-\bm{q}\cdot\bm{v}_{d,2}$ for the bottom layer.
@@
Eqs. (III)-(III) show that the vertical and lateral components of the current-modified Casimir interaction are composed of propagating ($q<\omega/c$) and evanescent ($q>\omega/c$) contributions, similar to the graphene-graphene interaction in equilibrium conditions [16]. We note that in the absence of a current, $F_{x}$ becomes zero as all the integrands are odd in $q_{x}$ leading to exact cancellations upon $q_{x}$ and $q_{x}$ integration. We also find that in this case, Eq. (III) can be evaluated by rotating the frequency integral into the complex plane, which transforms the $F_{z}$ expression into the conventional Lifhsitz-like form in the imaginary frequency domain [7, 15]. As a result, the Casimir force between the graphene sheets in the absence of drift currents is obtained as $F_{0}=-\frac{3\hbar c\alpha}{32\pi d^{4}}$ where $\alpha$ is the fine structure constant.
@@
At this point it is important to note that the Casimir interaction in the presence of drift currents has also been studied in a system composed of a nanoparticle and a 3D substrate, which is also relevant for the results obtained in Eqs. (III)-(III). In particular, Shapiro [26] developed a microscopic model for fluctuation-induced interactions between a nanoparticle and a current-carrying medium, identifying distinct physical pictures depending on whether the underlying fluctuations originate from the drifting electron plasma or from the lattice. In one of the considered scenarios, the contribution to the force comes from the moving electron plasma, while the lattice dissipation is neglected; in the other case, the fluctuation of the lattice dominates and thus the moving electrons would not affect the fluctuation-induced interaction. If the primary source of fluctuations originates from the lattice, it would be necessary to consider an effective temperature of the lattice that is different than the effective temperature of the drifting electrons, which complicates the problem as it is difficult to find these two different temperatures [26]. For the fluctuations originating primarily from the electrons a Doppler shift is induced in the fluctuation frequencies $\omega^{\prime},\omega^{\prime\prime}$ due to the carrier motion. In this work, we regard electrons as the primary source of fluctuations with a drift current in stationary medium.
Volokitin and Persson [30] modeled the drift current in one of the two parallel graphene layers via a Lorentz transformation—formally equivalent to a current-free medium moving along the $x$-axis at $v_{d}$. However, this treatment differs fundamentally from the physical situation in graphene, where the lattice remains static in the laboratory frame while only the carriers are moving. In the electron rest frame induced by the Lorentz boost, the lattice appears to move oppositely, yielding a dielectric response that deviates from the equilibrium case of static electrons and lattice. Therefore, a simple Doppler shift cannot adequately describe the graphene conductivity under drift. In [23] it has been shown that the SFD model and Doppler-shift approximation coincide only for an ideal two-dimensional electron gas, with discrepancies arising in graphene due to its Dirac nature. Here, we adopt a similar framework as in [30] while implementing the full SFD model to achieve a comprehensive description of the graphene optical response under drift.
@@ IV Current induced modifications in the interaction
To understand the role of drift currents in the graphene Casimir interaction, here we present results for the numerical calculations of Eqs. (III)-(III) for the case of charge carriers flowing along the $x$-axis in the top graphene layer with Doppler shifted frequency $\omega^{\prime}=\omega-q_{x}v_{\rm d}$ (for the bottom layer there is no drift velocity). The results in Fig. 3 (a)-(b) show how the modified interaction along the vertical direction depends on $\beta_{d}$ and $d$. At $\beta_{d}=0$ when there is no current, $\Delta F_{z}=0$ and we recover the usual equilibrium Casimir force between two graphene sheets $F_{0}=-\frac{3\hbar c\alpha}{32\pi d^{4}}$ previously discussed in literature [22, 7]. In the presence of a drift current, the Casimir force $F_{z}$ is modified by $\Delta F_{z}$ found to be with a positive sign. This repulsive graphene-graphene correction along the $z$-axis has a slightly slower than a quadratic $\beta_{d}^{2}$ dependence, as shown in Fig. 3 (a). We also note that $\Delta F_{z}$ has a different distance dependence when compared to $F_{0}$; in particular, we find that $\Delta F_{z}\sim\frac{1}{d^{5.8}}$ indicating a shorter range interaction compared to the equilibrium case with its $F_{0}\sim\frac{1}{d^{4}}$ distance power law. Fig. 3 (a, b) shows that this current-induced repulsive correction is most significant at small separations and large drift currents.
@@
As discussed earlier, due to the drifting electrons there is also a lateral Casimir-like force $F_{x}$ responsible for dissipating energy in the system. The lateral force opposes the direction of the driting carriers, and it is similar to the noncontact friction arising between two graphene sheets in relative motion [26, 8, 11] . Fig. 3 (c) shows that the lateral force is several orders of magnitude smaller when compared to the equilibrium $F_{0}$. It has a linear dependence upon $\beta_{d}$ for $\beta_{d}<1$, which is consistent with previously reported results for small drift current [30], and it has a nonlinear behavior as $\beta_{d}$ becomes larger. In Fig. 3 (d) we also show how $F_{x}$ changes as a function of the graphene-graphene separation. By interpolating the data, it is obtained that $F_{x}\sim\frac{1}{d^{4.2}}$ indicating that the lateral force has a somewhat shorter range coupling compared to $F_{0}$.
@@
It is also interesting to see how other factors, in addition to the interlayer distance and drift velocity, affect the lateral and vertical force components of the interaction. Fig. 4 (a) and (c) shows almost linear dependence upon temperature. The origin of this linear behavior can be traced to the thermal Bose factor entering the fluctuation-induced force. In the relevant low-frequency regime, where $\hbar\omega\ll k_{B}T$, the Bose–Einstein distribution admits the expansion $\left(\frac{1}{2}+n_{B}(\omega)\right)\approx\frac{k_{B}T}{\hbar\omega}+\mathcal{O}\left(\frac{\hbar\omega}{k_{B}T}\right)$, which shows that $\Delta F_{z}$ and $F_{x}$ become proportional to $k_{B}T/\hbar\omega$ to leading order. Another property that can also be used to modify the interaction is the graphene Fermi energy $E_{F}$. Increasing $E_{F}$ enhances the metallic-like nature of the materials leading to stronger Casimir interaction [7, 22, 10]. This is also observed in the case of current-modified graphene coupling. As shown in Fig. 4 (b) and (d), both the lateral and vertical forces experience direct correlation with $E_{F}$ with a close to linear functional behavior.
@@
At the end we discuss the case of currents flowing in both graphene layers but in the same direction. One notes that the two major factors in the Casimir energy in Eqs. (III)-(III) come from the Doppler shifted frequencies $\omega^{\prime}=\omega^{\prime\prime}=\omega-\bm{q}\cdot\bm{v}_{d}$ and the modified graphene conductivities according to the SFD model in Eq. (3). We can see that in this case the integrand in Eq. (III) is zero, which means that there is no lateral force. An analogous situation may also occur when the two graphene layers are physically moving with the same velocity in same directions [28, 14, 18]. The non-contact frictional force in this case is zero since there is no relative motion between electrons in the two layers. The correction to the vertical interaction $\Delta F_{z}$ is also small compared to the results in Figs.3 and 5. Since the Doppler shifted frequencies are not relevant due to absence of relative motion between the electrons of the two layers, only the SFD modified conductivities are important, therefore the corresponding correction to the equilibrium vertical component is also less prominent.